# Automatic Derivation of an Optimal Task Frame for Learning and Controlling Contact-Rich Tasks

### v1

@@ V-D Task Frame Derivation and Data Post-Processing
After transforming the data to the obtained task frame according to Section IV-E, the following signals are determined: $x,y,z$-positions and unit quaternions for pose, and the components of the twist and wrench screws. The signals were made time-invariant and aligned by reparameterizing them according to (15) as functions of the nondimensional geometric progress variable $\bar{\xi}$ that resulted from the task frame derivation, while using the same number of samples $N=100$ for all five trials. Next, we fitted a cubic spline to these aligned signals to find a smooth average over the trials. The quaternions are renormalized after smoothing.
@@ V-E Task Performance Setup and Control Strategy
The feedforward terms in (36) and (37) generate the exact required motion if the demonstration and robot scenarios are exactly the same. Hence, little or no contribution from the feedback terms is required in that case. If however the robot scenario deviates from the demonstration scenario, the feedback terms in the pose and wrench constraints come into play to compensate for these deviations. As a result, the desired control twists (36) and (37) are no longer consistent, but conflicting. To solve this conflict, weights $w_{p}$ and $w_{w}$ are applied to these equations, respectively, which results in a spring-damper like behavior, as depicted in Fig. 8. The equivalent stiffness and damping parameters, $k_{eq}$ and $b_{eq}$, are determined by gains $k_{p}$ and $k_{w}$, weights $w_{p}$ and $w_{w}$, and compliance $\underline{\Delta}\,\hat{\boldsymbol{C}}_{w}$. This relation is derived for a 1-d.o.f. system in Appendix B.
@@ VI-A Task Frame Derivation
Table IV summarizes the results of the task frame derivation. Columns 2-5 contain the results of the discrete decisions: viewpoints for the origin and the orientation, and the vectors of interest for motion and wrench. Each decision is based on the comparison of two relevant covariance matrices $\boldsymbol{C}_{1}$ and $\boldsymbol{C}_{2}$. Therefore, we also report a ratio,
@@
TABLE IV: Outcome of the proposed approach for selection of the task frame and comparison with task frames chosen by an expert
@@
Prismatic Joint $\mathcal{TL}\ (1.2)$ $\mathcal{TL}\ (1.3)$ $\boldsymbol{v}\ (1.3\text{e}^{1})$ $\boldsymbol{f}\ (1.6)$ $1.3^{\text{a}}$ $358.7^{\text{b}}$ $37.1^{\text{a}}$ $123^{\text{b}}$ $\mathbf{1.3}^{\text{a}}$ $\mathbf{166.5}^{\text{b}}$ $1.3^{\text{a}}$ $167.6^{\text{b}}$
Drawing $\mathcal{TL}\ (1.7\text{e}^{6})$ $\mathcal{W}\ (9.5)$ $\boldsymbol{v}\ (2.0)$ $\boldsymbol{f}\ (2.6)$ $0.3^{\text{c}}$ $498.8^{\text{d}}$ $3.9^{\text{c}}$ $8.9^{\text{d}}$ $3.8^{\text{c}}$ $\mathbf{8.9}^{\text{d}}$ $\mathbf{3.7}^{\text{c}}$ $\mathbf{8.9}^{\text{d}}$
@@
${}^{\text{a}}$ Total angle between $\prescript{[tf]}{}{\boldsymbol{e}}_{x}$ and $\prescript{[tf^{*}]}{}{\boldsymbol{e}}_{x}$              ${}^{\text{b}}$ Magnitude of common normal between $\prescript{[tf]}{}{\boldsymbol{e}}_{x}$ and $\prescript{[tf^{*}]}{}{\boldsymbol{e}}_{x}$
${}^{\text{c}}$ Total angle between $\prescript{[tf]}{}{\boldsymbol{e}}_{z}$ and $\prescript{[tf^{*}]}{}{\boldsymbol{e}}_{z}$              ${}^{\text{d}}$ $\left\lVert[p^{*}_{x}~{}p^{*}_{y}~{}p^{*}_{z}]^{T}-[p_{x}~{}p_{y}~{}p_{z}]^{T}\right\rVert$
${}^{\text{e}}$ $\left\lVert[p^{*}_{x}~{}p^{*}_{y}]^{T}-[p_{x}~{}p_{y}]^{T}\right\rVert$
Viewpoint. Viewpoint $\mathcal{TL}$ was consistently selected for the origin in all applications. The decision was particularly significant in all applications where the point of contact moved w.r.t. the world, for example in the Drawing task where the contact force is always applied at the tip of the tool, or which were demonstrated in different positions w.r.t. the world, for example in the Bottle Opening. Conversely, in applications such as Peg on Hole Alignment and Prismatic Joint, where the contact point does not move a lot w.r.t. the world, and/or which were always demonstrated in the same position w.r.t the world, the decision had a low significance. When analyzing the individual trials of the Peg on Hole Alignment we even observed that $\mathcal{TL}$ was favored for four trials while $\mathcal{W}$ was favored for one trial. Note that we can also consider the hole to be attached to the gripper and the peg to the world, which would result in swapping the viewpoints between $\mathcal{TL}$ and $\mathcal{W}$.
@@
Comparison with Expert’s Choice. Some, but certainly not all, of the task frame candidates derived from only motion or only wrench data closely align with the expert task frame. For example, the motion candidate for Revolute Joint and the wrench candidate for 2-D Contour Following closely align with $\{tf^{*}\}$, both in origin and orientation. Also, it appears that motion candidates are often more accurate, which is attributed to the operator exhibiting more variability in the wrench space. All this underscores the importance of averaging the results of motion and wrench candidates, which may involve combining complementary information. However, our results indicate that the weighted average of motion and wrench orientation candidates does not yield substantial benefit compared to the unweighted average, at least for the experiments that we performed.
@@ VI-B Control Performance
To assess the controller’s performance, the tracking errors for all signals and all robot experiments are reported in Table V as Root Mean Square Error (RMSE). The RMSE is defined as
$\Delta x=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left\lVert\boldsymbol{x}_{i}^{act}-\boldsymbol{x}_{i}^{des}\right\rVert^{2}},$ (39)
where $x$ is replaced by $p$, $\omega$, $v$, $f$ and $m$, while RMSE values for $\Delta R$ are calculated as $\left\lVert\boldsymbol{r}\right\rVert$, with $\boldsymbol{r}$ derived according to (4).
@@
Task No. $\Delta R$ $\Delta p$ $\Delta\omega$ $\Delta v$ $\Delta f$ $\Delta m$
@@ 3-D Contour Following
     Changed Origina 7-3 $3.8$ $26.6$ $1.7$ $2.8$ $9.4$ $0.2$
a Task performance was not successful.
@@ VI-B1 Nominal Scenarios
Figure 9: 2-D Contour Following performed by the UR10e robot: (a) nominal scenario with pose and wrench constraints; (b) nominal scenario without pose constraints.
@@
Independence of initial pose. This feature is illustrated in the nominal scenario of the Drawing task (5-1), see Fig. 10: the operator demonstrated using a tilted pen, where the initial angle between the normal vector of the surface and the pen was approximately 30, and approximately maintained a constant orientation during the task. In contrast, the robot started from an initial pose where the pen was approximately oriented along the normal of the surface, but was able to maintain a similar constant orientation w.r.t. this initial orientation. This is evidenced by the small value of $\Delta R$.
Figure 10: Difference in orientation of the tool in the Drawing task: (a) tilted in the demonstrations; (b) almost vertical in the robot performance
@@ VI-B2 Changed Geometry
Figure 11: Peg on Hole Alignment performed by the UR10e robot: (a) nominal scenario with peg diameter of 50 mm; (b) changed geometry with peg diameter of 35 mm.
Figure 12: Bottle Opening performed by the UR10e: (a) nominal scenario with same opener as in demonstrations; (b) changed geometry with a $50\%$ longer opener.
@@ VII Discussion and Conclusion
The main contribution of this article is an automatic procedure for deriving an optimal task frame for contact-rich robot tasks from demonstration data. The procedure is based on two principles that are hypothesized to underpin the control configuration targeted by an expert when choosing the task frame: (1) decoupled control of the directional and moment components of a screw, and (2) decoupled control of motion and interaction wrench. The procedure is rooted in screw theory and consists of a pipeline of algorithms. Besides existing algorithms such as ASIP and AVOF, the pipeline includes the following novelties: (1) an alternative application of ASIP (i.e. ASIP applied to a mean-subtracted screw), (2) a new algorithm to align orientation frames, (3) an adapted version of an algorithm to find the average of two orientations that can be applied to rotation matrices instead of quaternions, and (4) an extension of the concept of similarity transformation to pose matrices. The procedure is entirely probabilistic and does not involve any hyperparameters unless optional weighting is applied for averaging the orientation (but in our experiments we did not find evidence of significant added value for the latter). It also produces invariant results, regardless of the reference frames in which the data are expressed. It systematically uses (1) both motion and wrench data and (2) complete screw data, i.e. both in translation and orientation, and force and moment. This is innovative compared to the state of the art. The result is a gray-box model of the task, i.e. a set of parameters and reference signals that have unambiguous physical meaning and relations between them, and that are obtained in a data-driven way.
@@
The task frame derivation procedure was validated using seven applications including surface following tasks (2-D and 3-D Contour Following, Drawing), manipulation of articulated objects (Revolute and Prismatic joints), a nonreciprocal task involving plastic deformation (Bottle Opening) and tasks involving multiple contact points (Peg on Hole Alignment, 3-D Contour Following). The tasks involved both activities of daily life and technical tasks. While the derived task frames showed a good agreement with the task frames assumed to be chosen by experts, it is important to note that, (1) depending on the type of application, the accuracy of all the components of the origin or orientation may not be equally relevant, (2) in some cases the derived frame may be more optimal than the expert’s choice, which may be based on intuition or assumptions that are too simplified. Also, although the experiments were based on five demonstration trials for each application, the procedure can be applied to a single demonstration as well. Finally, our experiments confirmed the added value of combining motion and wrench data for deriving the optimal task frame.
@@
A second set of validation experiments consisted of a robot performing the seven tasks based on the obtained task model (task frame parameters and reference signals for twist, wrench and pose). We designed a one-size-fits-all controller based on the combination of pose and wrench constraints. Hence, the controller was certainly not optimal for all applications – that was not the main focus of the article – but could be tuned quickly for each application with only three parameters. The experiments showed that, for each application, successful task completion and stable control could be achieved in the nominal scenario where the robot performs the task under the same conditions as observed during the demonstrations. Furthermore, thanks to the type of controller and the gray-box nature of the task model, we could demonstrate successful task performances in case of (1) a different starting pose, (2) a changed task speed, (3) a changed magnitude of the interaction wrench and (4) small changes to the geometric configuration. For larger changes in the geometric configuration of tool or environment, we demonstrated two solutions: (1) manual adaptation of the task frame parameters and (2) iteration of the derivation of the task frame parameters using the data of the task performed by the robot. The latter option is interesting because such iteration eliminates calibration errors between the demonstration environment and the robot environment. Further investigation is required, but this option could potentially turn out to be a good standard practice after an application has been ported from demonstration to robot. Finally, it was shown that, due to its isotropic nature, the controller performed equally well regardless of the orientation of the task frame, while its performance deteriorated when the origin of the derived task frame was changed. The latter underscores the importance of selecting the optimal origin for the task frame.
@@
Obviously, the work reported in this paper has several limitations. First, we started from data that had already been segmented. Future work could investigate whether changes in the outcomes of the task frame derivation procedure can be used to identify transitions between the contact-rich segments studied in this article and other types of segments: gross motions in free space, approach-to-contact motions, building up of an initial contact and moving out of contact.
@@
Fifth, we derived an average task frame for the whole segment, with a fixed origin and a fixed orientation, both w.r.t. either world or tool. Although it was shown that such task frame is appropriate for many applications, an obvious next step for future work would be to pursue the most optimal task frame without these restrictions, i.e. an instantaneous task frame that is completely free to move. This would be beneficial for adequately controlling tasks with changing contact geometries, such as 2-D and 3-D Contour Following involving unknown objects or cutting.
@@
As a final remark, we believe that the task frame derivation approach fills a gap in the state of the art of LfD for contact-rich tasks, bringing this technology closer to practical applications while providing a solid base for further developments.

### v3

@@ V-D Task Frame Derivation and Data Post-Processing
After transforming the data to the obtained task frame according to Section IV-E, the following signals are determined: $x,y,z$-positions and unit quaternions for pose, and the components of the twist and wrench screws. The signals were made time-invariant and aligned to handle differences in velocity profile by reparameterizing them according to (15) as functions of the nondimensional geometric progress variable $\bar{\xi}$ that resulted from the task frame derivation, while using the same number of samples $N=100$ for all five trials. Next, we fitted a cubic spline to these aligned signals to find a smooth average over the trials. The quaternions are renormalized after smoothing.
@@ V-E Task Performance Setup and Control Strategy
The feedforward terms in (37) and (38) generate the exact required motion if the demonstration and robot scenarios are exactly the same. Hence, little or no contribution from the feedback terms is required in that case. If however the robot scenario deviates from the demonstration scenario, the feedback terms in the pose and wrench constraints come into play to compensate for these deviations. As a result, the desired control twists (37) and (38) are no longer consistent, but conflicting. To solve this conflict, weights $w_{p}$ and $w_{w}$ are applied to these equations, respectively, which results in a spring-damper like behavior, as depicted in Fig. 9. The equivalent stiffness and damping parameters, $k_{eq}$ and $b_{eq}$, are determined by gains $k_{p}$ and $k_{w}$, weights $w_{p}$ and $w_{w}$, and compliance $\underline{\Delta}\,\hat{\boldsymbol{C}}_{w}$. This relation can be shown for a 1-d.o.f. translational system,
$k_{eq}=\frac{w_{p}k_{p}}{w_{w}k_{w}\hat{C}_{f}}~{}\text{and}~{}~{}b_{eq}=\frac{1}{w_{w}k_{w}\hat{C}_{f}}$ (39)
where $\hat{C}_{f}$ is the translational compliance.
@@ VI-A Task Frame Derivation
Table IV summarizes the results of the task frame derivation. Columns 2-5 contain the results of the discrete decisions: viewpoints for the origin and the orientation ($\mathcal{W}$ or $\mathcal{TL}$), and the vectors of interest for motion and wrench ($\boldsymbol{v}$ or $\boldsymbol{\omega}$, and $\boldsymbol{f}$ or $\boldsymbol{m}$). Each decision is based on the comparison of two relevant covariance matrices $\boldsymbol{C}_{1}$ and $\boldsymbol{C}_{2}$. Therefore, between brackets, we also report a ratio,
@@
TABLE IV: Outcome of the proposed approach for selection of the task frame and comparison with task frames chosen by an expert.
The comparison with the expert for the Drawing task is visualized in Fig. 14. For the rest of the tasks, please refer to the videos in the supplementary material for the visualization.
@@
Prismatic Joint $\mathcal{TL}\ (1.2)$ $\mathcal{TL}\ (1.3)$ $\boldsymbol{v}\ (1.3\text{e}^{1})$ $\boldsymbol{f}\ (1.6)$ $1.3^{\text{a}}$ $358.7^{\text{b}}$ $37.1^{\text{a}}$ $122.9^{\text{b}}$ $\mathbf{1.3}^{\text{a}}$ $\mathbf{166.5}^{\text{b}}$ $1.3^{\text{a}}$ $167.6^{\text{b}}$
Drawing $\mathcal{TL}\ (1.7\text{e}^{3})$ $\mathcal{W}\ (9.5)$ $\boldsymbol{v}\ (2.0)$ $\boldsymbol{f}\ (2.6)$ $0.3^{\text{c}}$ $498.8^{\text{d}}$ $3.9^{\text{c}}$ $8.9^{\text{d}}$ $3.8^{\text{c}}$ $\mathbf{8.9}^{\text{d}}$ $\mathbf{3.7}^{\text{c}}$ $\mathbf{8.9}^{\text{d}}$
@@
${}^{\text{a}}$ Total angle between $\prescript{[tf]}{}{\boldsymbol{e}}_{x}$ and $\prescript{[tf^{*}]}{}{\boldsymbol{e}}_{x}$          ${}^{\text{b}}$ Magnitude of common normal between $\prescript{[tf]}{}{\boldsymbol{e}}_{x}$ and $\prescript{[tf^{*}]}{}{\boldsymbol{e}}_{x}$          ${}^{\text{c}}$ Total angle between $\prescript{[tf]}{}{\boldsymbol{e}}_{z}$ and $\prescript{[tf^{*}]}{}{\boldsymbol{e}}_{z}$
${}^{\text{d}}$ $\left\lVert[p^{*}_{x}~{}p^{*}_{y}~{}p^{*}_{z}]^{T}-[p_{x}~{}p_{y}~{}p_{z}]^{T}\right\rVert$                  ${}^{\text{e}}$ $\left\lVert[p^{*}_{x}~{}p^{*}_{y}]^{T}-[p_{x}~{}p_{y}]^{T}\right\rVert$
Viewpoint. Viewpoint $\mathcal{TL}$ was consistently selected for the origin in all applications. The decision was particularly significant in all applications where the point of contact moved w.r.t. the world, for example in the Drawing task where the contact force is always applied at the tip of the tool, or which were demonstrated in different positions w.r.t. the world, for example in the Bottle Opening. Conversely, in applications such as Peg on Hole Alignment and Prismatic Joint, where the contact point does not move a lot w.r.t. the world, and/or which were always demonstrated in the same position w.r.t. the world, the decision had a low significance. When analyzing the individual trials of the Peg on Hole Alignment we even observed that $\mathcal{TL}$ was favored for four trials while $\mathcal{W}$ was favored for one trial. Note that we can also consider the hole to be attached to the gripper and the peg to the world, which would result in swapping the viewpoints between $\mathcal{TL}$ and $\mathcal{W}$.
@@
Comparison with Expert’s Choice. Some, but certainly not all, of the task frame candidates derived from only motion or only wrench data closely align with the expert task frame. For example, the motion candidate for Revolute Joint and the wrench candidate for 2-D Contour Following closely align with $\{tf^{*}\}$, both in origin and orientation. Also, it appears that motion candidates are often more accurate, which is attributed to the operator exhibiting more variability in the wrench space. All this underscores the importance of averaging the results of motion and wrench candidates, which may involve combining complementary information. The comparison with the expert for the Drawing task is visualized in Fig. 14. In this application, the origin of the task frame is attached to the tool while the orientation of the task frame is attached to the world. For the rest of the tasks, please refer to the videos in the supplementary material for the visualization of the derived and expert task frames. Finally, our results indicate that the weighted average of motion and wrench orientation candidates does not yield substantial benefit compared to the unweighted average, at least for the experiments that we performed.
@@ VI-B Control Performance
To assess the controller’s performance, the tracking errors for all signals and all robot experiments are reported in Table V using the RMSE.
@@
Task No. $\Delta R^{a}$ $\Delta p$ $\Delta\omega$ $\Delta v$ $\Delta f$ $\Delta m$
@@ 3-D Contour Following
     Changed Originb 7-3 $3.8$ $26.6$ $1.7$ $2.8$ $9.4$ $0.2$
a The RMSE values for $\Delta R$ are calculated as $\left\lVert\boldsymbol{r}\right\rVert$, with $\boldsymbol{r}$ derived according to (4).
b Task performance was not successful.
@@ VI-B1 Nominal Scenarios
Figure 10: 2-D Contour Following performed by the UR10e robot (snapshots of video): (a) nominal scenario with pose and wrench constraints; (b) nominal scenario without pose constraints, resulting in task failure.
@@
Independence of initial pose. This feature is illustrated in the nominal scenario of the Drawing task (5-1), see Fig. 11b: the operator demonstrated using a tilted pen, where the initial angle between the normal vector of the surface and the pen was approximately 30, and approximately maintained a constant orientation during the task. In contrast, the robot started from an initial pose where the pen was approximately oriented along the normal of the surface, but was able to maintain a similar constant orientation w.r.t. this initial orientation. This is evidenced by the small value of $\Delta R$.
Figure 11: Difference in orientation of the tool in the Drawing task (snapshots of video): (a) tilted in the demonstrations; (b) almost vertical in the robot performance
@@ VI-B2 Changed Geometry
Figure 12: Peg on Hole Alignment performed by the UR10e robot (snapshots of video): (a) nominal scenario with peg diameter of 50 mm; (b) changed geometry with peg diameter of 35 mm.
Figure 13: Bottle Opening performed by the UR10e (snapshots of video): (a) nominal scenario with same opener as in demonstrations; (b) changed geometry with a $50\%$ longer opener.
@@ VI-B3 Sensitivity to Task Frame Origin and Orientation
Figure 14: Expert task frame (pale) and derived task frame (bright) for the Drawing task according to the RGB convention. The origin of the task frame is attached to the tool while the orientation of the task frame is attached to the world. The derived task frames for the other applications are shown in videos in the supplementary material.
@@ VII Discussion and Conclusion
Contributions. The main contribution of this article is an automatic procedure for deriving an optimal task frame for contact-rich robot tasks from demonstration data. The procedure is summarized in Fig. 6 and fits in the red block of the overall workflow depicted in Fig. 1b. The demonstration data may consist of human-demonstrated, simulation or robot data. The procedure is based on two principles that are hypothesized to underpin the control configuration targeted by an expert when choosing the task frame: (1) decoupled control of the directional and moment components of a screw, and (2) decoupled control of motion and interaction wrench. The procedure is rooted in screw theory and consists of a pipeline of algorithms. Besides existing algorithms such as ASIP and AVOF, the pipeline includes the following novelties: (1) an alternative application of ASIP (i.e. ASIP applied to a mean-subtracted screw), (2) a new algorithm to align orientation frames, (3) an adapted version of an algorithm to find the average of two orientations that can be applied to rotation matrices instead of quaternions, and (4) an extension of the concept of similarity transformation to pose matrices. The procedure is entirely probabilistic and does not involve any hyperparameters unless optional weighting is applied for averaging the orientation (but in our experiments we did not find evidence of significant added value for the latter). It also produces invariant results, regardless of the reference frames in which the data are expressed. The result is a gray-box model of the task, i.e. a set of parameters and reference signals that have unambiguous physical meaning and relations between them, and that are obtained in a data-driven way.
Main innovations with respect to the state of the art. The task frame derivation approach is fully automatic and fully data-driven. I.e., it is generative, as it does not require predefined candidate reference frames or any input other than the demonstration data. Furthermore, the task frame is not constrained to be entirely fixed to the world or the tool as the origin and the orientation may be - independently - fixed to either world or tool. The approach systematically uses (1) both motion and wrench data and (2) complete screw data, i.e. both in translation and orientation, and force and moment. As a result, the approach is generic, as it applies to a large variety of contact-rich tasks.
@@
Validation of the task frame derivation procedure. First, the task frame derivation procedure itself was validated using seven applications including surface following tasks (2-D and 3-D Contour Following, Drawing), manipulation of articulated objects (Revolute and Prismatic joints), a nonreciprocal task involving plastic deformation (Bottle Opening) and tasks involving multiple contact points (Peg on Hole Alignment, 3-D Contour Following). The tasks involved both activities of daily life and technical tasks. While the derived task frames showed a good agreement with the task frames assumed to be chosen by experts, it is important to note that, (1) depending on the type of application, the accuracy of all the components of the origin or orientation may not be equally relevant, (2) in some cases the derived frame may be more optimal than the expert’s choice, which may be based on intuition or assumptions that are too simplified. Also, although the experiments were based on five demonstration trials for each application, the procedure can be applied to a single demonstration as well. Finally, our experiments confirmed the added value of combining motion and wrench data for deriving the optimal task frame.
@@
Validation of the benefits of using a well-chosen and parameterized task frame model. A second set of validation experiments consisted of a robot performing the seven tasks based on the obtained task model (task frame parameters and reference signals for twist, wrench and pose). We designed a one-size-fits-all controller based on the combination of pose and wrench constraints. Hence, the controller was certainly not optimal for all applications – that was not the main focus of the article – but could be tuned quickly for each application with only three parameters. The experiments showed that, for each application, successful task completion and stable control could be achieved in the nominal scenario where the robot performs the task under the same conditions as observed during the demonstrations. Furthermore, thanks to the type of controller and the gray-box nature of the task model, we could demonstrate successful task performances in case of (1) a different starting pose, (2) a changed task speed, (3) a changed magnitude of the interaction wrench and (4) small changes to the geometric configuration. For larger changes in the geometric configuration of tool or environment, we demonstrated two solutions: (1) manual adaptation of the task frame parameters and (2) iteration of the derivation of the task frame parameters using the data of the task performed by the robot. The latter option is interesting because such iteration eliminates calibration errors between the demonstration environment and the robot environment. Further investigation is required, but this option could potentially turn out to be a good standard practice after an application has been ported from demonstration to robot. Finally, it was shown that, due to its isotropic nature, the controller performed equally well regardless of the orientation of the task frame, while its performance deteriorated when the origin of the derived task frame was changed. The latter underscores the importance of selecting the optimal origin for the task frame.
@@
Limitations and future work. Current limitations of the proposed method are the following. First, we started from data that had already been segmented. Future work could investigate whether changes in the outcomes of the task frame derivation procedure can be used to identify transitions between the contact-rich segments studied in this article and other types of segments: gross motions in free space, approach-to-contact motions, building up of an initial contact and moving out of contact.
@@
Fifth, we derived an average task frame for the whole segment, with a fixed origin and a fixed orientation, both w.r.t. either world or tool. A future extension would be to derive an optimal task frame without these restrictions, i.e. an instantaneous task frame that is completely free to move. Such frame would be beneficial for controlling tasks with changing contact geometries, as in 2-D and 3-D Contour Following involving unknown objects or cutting. This could be achieved by considering the average task frame as proposed in this paper, but over a shorter time window, or by switching to invariant motion and wrench trajectory descriptors as in [41].